# Unleashing the Power of Task-Specific Directions in Parameter Efficient Fine-tuning

### v2

@@ 3.4 Challenges and Explorations: Using TSD is Not as Easy as Expected
Despite the apparent challenges, we remain undeterred and placed our hopes on $\Delta\mathbf{W}$. We hypothesized that the core directions with the highest change rates predicted by $\Delta\mathbf{W}$ from LoRA are strongly associated with TSDs. To test this hypothesis, we conducted extensive experiments. We strongly encourage readers to refer to Sec. D.3 for detailed experimental setups. The results are shown in the Figs. 2 and 7, where we observe a significant overlap between the predicted directions and the actual TSD, highlighting their strong correlation. It leads to an important conclusion:
Proposition 1.
@@ 5.1 Numerical Results
Robustness to Parameter Budget: (When fine-tuning LLaMA-7B) LoRA shows enhanced performance with an increased parameter budget, indicating its dependence on larger rank sizes. This highlights its effectiveness in harnessing the full potential of TSDs, even in constrained settings. Conversely, LoRA-Dash maintains strong performance even under limited parameter conditions, which suggests that TSDs are crucial for maximizing fine-tuning efficiency.
Surpassing Fully Fine-Tuning Performance: While LoRA typically falls short of the results achieved by fully fine-tuning, LoRA-Dash in some cases outperforms full fine-tuning. This intriguing observation may provide new insights into optimizing training strategies for fine-tuning.
For comparisons with other methods, please refer to Tables. 22-23.
@@ 5.2 Visual Results
Moreover, the effectiveness of TSDs is particularly evident in the response of LoRA-Dash to given prompts. TSDs effectively encapsulate the specific information of each prompt, enabling LoRA-Dash to precisely grasp the semantic essence of prompts such as sea beach or wheat field. This capability allows LoRA-Dash to adeptly render complex themes and nuanced details within the images.
@@ 6.1 How Does LoRA-Dash Determine and Align with Task-specific Directions
The first is to explore the consistency between the directions selected during the pre-launch phase of LoRA-Dash and those showing the greatest change after the final training, and the second and third aim to validate the consistency between the directions chosen during the LoRA-Dash pre-launch phase and the final trained with the true TSDs. We conducted tests on LLaMA-7B Touvron et al. (2023a) and DeBERTaV3-large He et al. (2021b), each with three ranks of LoRA-Dash, and the results are shown in Tables. 3 and 20-21.
Table 3: Alignment on the directions of LoRA-Dash and TSDs. We report the percentage from the self-attention value (v) projection of the first, middle and last layer of two models, as well as the average percentage of all q, k, and v modules. More results are shown in Tables. 20-21.
@@ D.1 Implementation Details
The results presented in Figs 2 and 7 lead to two significant observations regarding the effectiveness of using $\Delta\mathbf{W}$ to identify TSDs:
@@ E.2 More Results for the Alignment on the Directions of LoRA-Dash
More results for Sec. 6.1 are shown in Tables. 20-21.
@@ E.6 Comparison with Other Methods
Specifically, we compare BitFit Zaken et al. (2021), Prompt Learning (Prefix) Li & Liang (2021), Series Adapter (Series) Houlsby et al. (2019), Adapter proposed by Pfeiffer et al. (2020) (PAdapter), Parallel Adapter (Parallel) He et al. (2021a), (IA)3 Liu et al. (2022), SSL Si et al. (2024b), SSB Si et al. (2024b), AdaLoRA Zhang et al. (2022), PISSA Meng et al. (2024), MiLoRA Wang et al. (2024), FLoRA Si et al. (2024a) and DoRA Liu et al. (2024). We follow the settings detailed in our paper, and the results are shown in Tables. 22-23.

### v3

@@ 3.4 Challenges and Explorations: Using TSD is Not as Easy as Expected
Despite the apparent challenges, we place our confidence in the potential of $\Delta\mathbf{W}$. We hypothesized that the core directions with the highest change rates predicted by $\Delta\mathbf{W}$ from LoRA are strongly associated with TSDs. To test this hypothesis, we conducted extensive experiments. We strongly encourage readers to refer to Sec. D.3 for detailed experimental setups. The results are shown in the Figs. 2 and 8, where we observe a significant overlap between the predicted directions and the actual TSD, highlighting their strong correlation. It leads to an important conclusion:
Observation 1.
@@ 5.1 Numerical Results
Robustness to Parameter Budget: (When fine-tuning LLaMA-7B) LoRA shows enhanced performance with an increased parameter budget, indicating its dependence on larger rank sizes. Conversely, LoRA-Dash maintains strong performance even under limited parameter conditions, which suggests that TSDs are crucial for maximizing fine-tuning efficiency. This highlights its effectiveness in harnessing the full potential of TSDs, even in constrained settings.
Surpassing Fully Fine-Tuning Performance: While LoRA typically falls short of the results achieved by fully fine-tuning, LoRA-Dash in some cases outperforms fully fine-tuning. This intriguing observation may provide new insights into optimizing training strategies for fine-tuning.
For comparisons with other methods, please refer to Tables. 23-24.
@@ 5.2 Visual Results
Moreover, the effectiveness of TSDs is particularly evident in the response of LoRA-Dash to given prompts. TSDs effectively encapsulate the specific information of each prompt, enabling LoRA-Dash to precisely grasp the semantic essence of prompts such as sea beach or wheat field. This capability allows LoRA-Dash to adeptly render complex themes and details within the images.
@@ 6.1 How Does LoRA-Dash Determine and Align with Task-specific Directions
The first is to explore the consistency between the directions selected during the pre-launch phase of LoRA-Dash and those showing the greatest change after the final training, and the second and third aim to validate the consistency between the directions chosen during the LoRA-Dash pre-launch phase and the final trained with the true TSDs. We conducted tests on LLaMA-7B Touvron et al. (2023a) and DeBERTaV3-large He et al. (2021b), each with three ranks of LoRA-Dash, and the results are shown in Tables. 3 and 21-22.
Table 3: Alignment on the directions of LoRA-Dash and TSDs. We report the percentage from the self-attention value (v) projection of the first, middle and last layer of two models, as well as the average percentage of all q, k, and v modules. More results are shown in Tables. 21-22.
@@ 7 Conclusion
Acknowledgements
This work was supported by NSFC 62322604, NSFC 62176159, and Shanghai Municipal Science and Technology Major Project 2021SHZDZX0102.
@@ D.1 Implementation Details
Additionally, it is worth noting that the phenomenon presented in Fig. 1 is not an isolated case. Similar behavior has been observed in other tasks and models. For instance, as shown in Fig., results obtained by training Qwen2.5-7B Team (2024) on the math reasoning task Hu et al. (2023) exhibit patterns that are fundamentally consistent with our observations in Fig. 1.
Figure 7: The results when fine-tuning Qwen2.5-7B on math reasoning task are similar to those in Fig. 1
@@
The results presented in Figs 2 and 8 lead to two significant observations regarding the effectiveness of using $\Delta\mathbf{W}$ to identify TSDs:
@@ E.1 More Results for Subject-driven Generation Task
We also evaluated the performance using the text-image score evaluated by CLIP Radford et al. (2021) and the image similarity score. The CLIP text-image score measures the alignment between textual descriptions and the corresponding images, while the image similarity score quantifies the similarity between the original image and the generated image. The results are presented in Table 20, where it is evident that LoRA-Dash outperforms LoRA on both metrics. The images generated by LoRA-Dash exhibit better text alignment and a stronger correspondence with the original images, demonstrating that LoRA-Dash achieves superior results compared to LoRA.
Table 20: Performance comparison between LoRA and LoRA-Dash on CLIP text-image alignment and image similarity metrics.
Metric (%) LoRA LoRA-Dash
Mean CLIP Text-Image Score 32.0 32.8
Mean Image Similarity 76.4 78.2
@@ E.2 More Results for the Alignment on the Directions of LoRA-Dash
More results for Sec. 6.1 are shown in Tables. 21-22.
@@ E.6 Comparison with Other Methods
Specifically, we compare BitFit Zaken et al. (2021), Prompt Learning (Prefix) Li & Liang (2021), Series Adapter (Series) Houlsby et al. (2019), Adapter proposed by Pfeiffer et al. (2020) (PAdapter), Parallel Adapter (Parallel) He et al. (2021a), (IA)3 Liu et al. (2022), SSL Si et al. (2024b), SSB Si et al. (2024b), AdaLoRA Zhang et al. (2022), PISSA Meng et al. (2024), MiLoRA Wang et al. (2024), FLoRA Si et al. (2024a) and DoRA Liu et al. (2024). We follow the settings detailed in our paper, and the results are shown in Tables. 23-24.